### v2

@@ 6 Conclusion
One limitation is that there is still a lot of room for improvement in the accuracy of our model in Recognizing hand-drawn molecules and molecules in mechanism images as we mentioned before. Another is that in real literature, R-group information in molecular images often appears in other places, such as texts or tables [41, 42]. future work may need to integrate these information to obtain a more complete molecular prediction.
@@
List of abbreviations OCSR: Optical Chemical Structure Recognition; SMILES: Simplifed Molecular-Input Line-entry System; SMARTS: SMiles ARbitrary Target Specification; CNN: Convolutional Neural Network; ViT: Vision-Transformer; OCR:Optical Character Recognition; RNN: Recurrent Neural Networks; FFN:Feed-Forward Network; MHSA: Multi-Head Self-Attention
@@
Funding HKUST (Project No. R9251, Z1269)

### v3

@@ 6 Conclusions
One limitation is that there is still a lot of room for improvement in the accuracy of our model in Recognizing hand-drawn molecules and molecules in mechanism images as we mentioned before. Another is that in real literature, R-group information in molecular images often appears in other places, such as texts or tables [41, 42]. Future work may need to integrate this information to obtain a more complete molecular prediction.
@@
List of abbreviations OCSR: Optical Chemical Structure Recognition; MolNexTR: Molecular convNext-TRansformer; SMILES: Simplifed Molecular-Input Line-entry System; SMARTS: SMiles ARbitrary Target Specification; CNN: Convolutional Neural Network; ViT: Vision-Transformer; OCR: Optical Character Recognition; RNN: Recurrent Neural Networks; FFN: Feed-Forward Network; MHSA: Multi-Head Self-Attention; GPUs: Graphical Processing Units; Render Aug.: Render Augmentation; Image Aug.: Image Augmentation; Mol Aug.: Molecular Augmentation; Img Contam Alg.: Image Contamination Algorithm.
@@
Funding This work was supported by HKUST (Project No. R9251, Z1269)